# Banach algebras associated to twisted étale groupoids: simplicity and pure infiniteness

### v2

@@ Acknowledgements
The research of KB and BK was supported by the National Science Center (NCN) grant no. 2019/35/B/ST1/02684. BK was partially supported by the National Science Centre, Poland, through the WEAVE-UNISONO grant no. 2023/05/Y/ST1/00046. BK thanks Agnieszka Stocka whose comments helped formulate the definition of reduced algebras in a correct way. We all thank anonymous referees for pointing us to [Dix61] and [Elk24].
@@ 2. Preliminaries
We recall here some notation and facts from [BKM23]. Throughout this paper $\mathcal{G}$ denotes an étale groupoid with locally compact Hausdorff unit space $X$. Thus the range and domain maps $r,d:\mathcal{G}\to X$ are local homeomorphisms, $X$ is an open subset of $\mathcal{G}$, and $\mathcal{G}$ is Hausdorff if and only if $X$ is also closed in $\mathcal{G}$. We consider spaces, and algebras, over the field $\mathbb{F}=\mathbb{R},\mathbb{C}$ of real or complex numbers. By an ideal in a Banach algebra, unless stated otherwise, we mean a closed and two-sided ideal.
@@ 2.1. Twisted groupoids
As in [BKM23, Section 4], we define twists of $\mathcal{G}$ as Fell line bundles. A line bundle $\mathcal{L}$ over $\mathcal{G}$ is a topological space $\mathcal{L}=\bigsqcup_{\gamma\in\mathcal{G}}L_{\gamma}$ such that the canonical map $\mathcal{L}\to\mathcal{G}$ is continuous and open,  and each fiber $L_{\gamma}\cong\mathbb{F}$ is a one dimensional Banach space with a structure consistent with the topology of $\mathcal{L}$ in the sense that the maps $\bigsqcup_{\gamma\in\mathcal{G}}L_{\gamma}\times L_{\gamma}\ni(z,w)\mapsto z+w\in\mathcal{L}$, $\mathbb{F}\times\mathcal{L}\ni(\lambda,z)\mapsto\lambda z\in\mathcal{L}$ and $\mathcal{L}\ni z\mapsto|z|\in\mathbb{R}$ are continuous. For any open $U\subseteq\mathcal{G}$ we write $C_{c}(U,\mathcal{L})$, $C_{0}(U,\mathcal{L})$, $C(U,\mathcal{L})$ and $C_{u}(U,\mathcal{L})$ for respectively the spaces of continuous sections of $\mathcal{L}|_{U}:=\bigsqcup_{\gamma\in U}L_{\gamma}$ that are compactly supported, vanishing at infinity, bounded and unitary, respectively. We use similar notation $\mathcal{B}_{c}(U,\mathcal{L})$, $\mathcal{B}_{0}(U,\mathcal{L})$, $\mathcal{B}(U,\mathcal{L})$ when continuous is replaced by Borel. In particular, $\mathcal{B}_{0}(U,\mathcal{L})$ consists of Borel sections $f:U\to\mathcal{L}$ such that for every $\varepsilon>0$ the set $\{\gamma\in U:|f(\gamma)|\geq\varepsilon\}$ is precompact. We consider these spaces equipped with the supremum norm $\|f\|_{\infty}=\sup_{\gamma\in U}|f(\gamma)|$. Every line bundle is locally trivial which means that every point has a neighborhood $U\subseteq\mathcal{G}$ such that $C_{u}(U,\mathcal{L})\neq\emptyset$: for any $c\in C_{u}(U,\mathcal{L})$ the map $f\mapsto f\cdot c$ establishes the isometric isomorphism $C_{c}(U,\mathcal{L})\cong C_{c}(U)$ that extends to isomorphims between the other aforementioned spaces.
@@
A twist over the groupoid $\mathcal{G}$ is a line bundle $\mathcal{L}=\bigsqcup_{\gamma\in\mathcal{G}}L_{\gamma}$ over $\mathcal{G}$ equipped with two continuous maps :×$\cdot:\mathcal{L}\times\mathcal{L}\supseteq\bigsqcup_{d(\gamma)=r(\eta)}L_{\gamma}\times L_{\eta}\to\mathcal{L}$ and ${}^{*}:\mathcal{L}\to\mathcal{L}$ called multiplication and involution, such that:
@@
    $\cdot$ restricts to bilinear maps $L_{\gamma}\times L_{\eta}\to L_{\gamma\eta}$, $(\gamma,\eta)\in\mathcal{G}^{(2)}$, that make the norm ||$\cdot$|| ⋅ | multiplicative, and $\cdot$ is associative, i.e. $(\mathcal{L},\cdot)$ is a semigroupoid;
@@
The trivial bundle $\mathcal{L}=\mathcal{G}\times\mathbb{F}$ with pointwise operations is the trivial twist. More generally, if $\sigma:\mathcal{G}^{(2)}\to\{z\in\mathbb{F}:|z|=1\}$ is a continuous normalised $2$-cocycle on $\mathcal{G}$, i.e. $\sigma\big{(}r(\gamma),\gamma\big{)}=1=\sigma\big{(}\gamma,d(\gamma)\big{)}$ and $\sigma(\alpha,\beta)\sigma(\alpha\beta,\gamma)=\sigma(\beta,\gamma)\sigma(\alpha,\beta\gamma)$ for every composable triple $\alpha,\beta,\gamma\in\mathcal{G}$, then we may consider a twist $\mathcal{L}_{\sigma}$ given by the trivial bundle $\mathcal{G}\times\mathbb{F}$ with operations given by $(\alpha,w)\cdot(\beta,z):=\big{(}\alpha\beta,\sigma(\alpha,\beta)wz\big{)}$, $(\alpha,w)^{*}:=\big{(}\alpha^{-1},\overline{\sigma(\alpha^{-1},\alpha)w}\big{)}.$ Every twist based on a trivial line bundle is of this form. In general, twists defined above are equivalent to Kumjian’s twist [Kum86] defined using extensions of $\mathcal{G}$.
The set of open bisections $\operatorname{Bis}(\mathcal{G}):=\{U\subseteq\mathcal{G}\text{ open}:\text{$r|_{U}$ and $d|_{U}$ are injective}\}$, with operations induced from $\mathcal{G}$, that is for $U,V\in\operatorname{Bis}(\mathcal{G})$ we have
@@
forms a unital inverse semgroup where the unit is $X$. For any twist $\mathcal{L}$ over $\mathcal{G}$ the set
$S(\mathcal{L}):=\{U\in\operatorname{Bis}(\mathcal{G}):\text{the restricted bundle $\mathcal{L}|_{U}$ is trivial}\}$
forms a unital inverse subsemigroup $S(\mathcal{L})\subseteq\operatorname{Bis}(\mathcal{G})$ which is wide in the sense that it covers $\mathcal{G}$ and the intersection of any two elements in $S(\mathcal{L})$ is a union of elements from $S(\mathcal{L})$, see [BKM23, Lemma 4.2].
@@ 2.2. Banach algebras associated to twisted groupoids
where $\cdot$ and the final * indicate the product and involution from $\mathcal{L}$. When the groupoid $\mathcal{G}$ is Hausdorff this reduces to the usual *-algebra $C_{c}(\mathcal{G},\mathcal{L})$. We associate a norm on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ with any unital inverse subsemigroup $S\subseteq\operatorname{Bis}(\mathcal{G})$ which covers $\mathcal{G}$. By [BKM23, Lemma 4.5], the formula
@@
The class of algebras $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ defined above consists of Banach algebra Hausdorff completions of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ in seminorms $\|\cdot\|_{\mathcal{R}}$ for which there exists a unital inverse subsemigroup $S\subseteq\operatorname{Bis}(\mathcal{G})$ covering $\mathcal{G}$ such that $\|\cdot\|_{\mathcal{R}}\leq\|\cdot\|_{\infty}$ on every $C_{c}(U,\mathcal{L})$, $U\in S$ (this is equivalent to $\|\cdot\|_{\mathcal{R}}\leq\|\cdot\|_{\textrm{max}}^{S}$ on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$). Without loss of generality, we may always assume here that $S$ is closed under taking subsets, so that in particular it is wide, and that $S\subseteq S(\mathcal{L})$, see [BKM23, Remark 4.7].
@@
Let $p\in[1,\infty]$. We recall (see e.g. [Phi13b]) that a Banach algebra $A$ is an $L^{p}$-operator algebra if there is an isometric representation $\psi:A\to B(L^{p}(\mu))$, where $L^{p}(\mu)$ is the Banach $L^{p}$-space associated to some measure $\mu$. Following [BKM23, Definition 5.9] we define the full $L^{p}$-operator algebra of $(\mathcal{G},\mathcal{L})$ as $F^{p}(\mathcal{G},\mathcal{L}):=F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ where $\mathcal{R}$ consists of representations $\psi:F^{S}(\mathcal{G},\mathcal{L})\to B(E)$ for some unital inverse subsemigroup $S\subseteq\operatorname{Bis}(\mathcal{G})$ which covers $\mathcal{G}$, such that
@@
    $\psi|_{C_{0}(X)}$ acts by multiplication operators.
The following comments hold by [BKM23, Theorem 5.10]. The above definition does not depend on the choice of $S$, and so $F^{p}(\mathcal{G},\mathcal{L})$ is $\operatorname{Bis}(\mathcal{G})$-graded. In fact denoting by $\|\cdot\|_{L^{p}}$ the norm in $F^{p}(\mathcal{G},\mathcal{L})$, for $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ and $q\in[1,\infty]$ with $11/p+1/q=$ 1 (we put $q=1$ if $p=\infty$ and vice versa) we get
@@
in the complex case. We do not know whether this holds in the real case, and a potential trouble may come from representations as in [BKM23, Example 2.13]. For $\mathbb{F}=\mathbb{R},\mathbb{C}$ and $p=2$, $\|\cdot\|_{L^{2}}$ is the maximal submultiplicative norm on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ satisfying the $C^{*}$-equality. Thus $C^{*}(\mathcal{G},\mathcal{L}):=F^{2}(\mathcal{G},\mathcal{L})$ is a (real or complex) $C^{*}$-algebra, which is universal in the sense that *-homomorphisms from $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ into a $C^{*}$-algebra $B$ extend to representations of $C^{*}(\mathcal{G},\mathcal{L})$ in $B$, and every representation of $C^{*}(\mathcal{G},\mathcal{L})$ in $B$ is automatically *-preserving. For $p=\infty$ condition (R1) can be replaced by: $E=C_{0}(\Omega)$ for some locally compact Hausdorff space. Thus $F^{\infty}(\mathcal{G},\mathcal{L})$ could also be viewed as the full $C_{0}$-operator algebra of $(\mathcal{G},\mathcal{L})$. Moreover, if $q\in[1,\infty]$ satisfies $11/p+1/q=$ 1, then the involution in $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ extends to an isometric anti-isomorphism $F^{p}(\mathcal{G},\mathcal{L})\stackrel{{\scriptstyle\mathrm{anti}}}{{\cong}}F^{q}(\mathcal{G},\mathcal{L})$.
@@
defines a representation $\Lambda_{p}:F^{p}(\mathcal{G},\mathcal{L})\to B(\ell^{p}(\mathcal{G},\mathcal{L}))$ which is injective on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, see [BKM23, Proposition 5.1]. We call $F_{\operatorname{r}}^{p}(\mathcal{G},\mathcal{L}):=F_{\operatorname{r}}^{\{\Lambda_{p}\}}(\mathcal{G},\mathcal{L})$ the reduced $L^{p}$-operator algebra of $(\mathcal{G},\mathcal{L})$ and we denote the norm by $\|\cdot\|_{L^{p},\operatorname{r}}$. By [BKM23, Theorem 5.10] we have
@@ 3. Reduced groupoid Banach algebras
We fix a twisted groupoid $(\mathcal{G},\mathcal{L})$. In this section we will consider completions of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ that are equipped with ‘generalized conditional expectations’ that extend the map $E_{X}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to\mathcal{B}_{c}(X)$, given by restriction $f\mapsto f|_{X}$. When $\mathcal{G}$ is Hausdorff, then $E_{X}:C_{c}(\mathcal{G},\mathcal{L})\to C_{c}(X)\subseteq C_{c}(\mathcal{G},\mathcal{L})$ is a projection.
@@
For any unital inverse subsemigroup $S\subseteq\operatorname{Bis}(\mathcal{G})$ covering $\mathcal{G}$, the algebra $F^{S}(\mathcal{G},\mathcal{L})$ is a groupoid Banach algebra. For any $p\in[1,\infty]$ the $L^{p}$-operator algebras $F^{p}(\mathcal{G},\mathcal{L})$ and $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ are groupoid Banach algebras and $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ is reduced, cf. Proposition 3.15 below. There is a unique $C^{*}$-norm $\|\cdot\|_{C^{*}_{\operatorname{r}}}$ on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ yielding a reduced Banach algebra, and then this algebra coincides with the standard reduced $C^{*}$-algebra $C^{*}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ of $(\mathcal{G},\mathcal{L})$ (or its real counterpart). The full $C^{*}$-algebra $C^{*}(\mathcal{G},\mathcal{L})$, which is a completion of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ in the maximal $C^{*}$-norm $\|\cdot\|_{C^{*}\max}$, is a groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$. Thus a $C^{*}$-norm $\|\cdot\|_{\mathcal{R}}$ on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ yields a groupoid Banach algebra if and only if it satisfies $\|\cdot\|_{C^{*}_{\operatorname{r}}}\leq\|\cdot\|_{\mathcal{R}}\leq\|\cdot\|_{C^{*}\max}$.
We make some preparations to reveal more structure in groupoid Banach algebras. For any $U\in\operatorname{Bis}(\mathcal{G})$ we view the space of bounded Borel sections $\mathcal{B}(U,\mathcal{L})$ as a Banach $C(X)$-bimodule:
@@
gives a $C_{0}(X)$-bilinear map $E_{U}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to\mathcal{B}(U,\mathcal{L})$. The map $E_{U}$ on $\mathfrak{C}_{c}(r(U)\mathcal{G},\mathcal{L})=C_{c}(r(U))*\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, takes values in compactly supported Borel sections $\mathcal{B}_{c}(U,\mathcal{L})=C_{c}(r(U))*\mathcal{B}(U,\mathcal{L})$. When $\mathcal{G}$ is Hausdorff, then in fact $E_{U}:\mathfrak{C}_{c}(r(U)\mathcal{G},\mathcal{L})\to C_{c}(U,\mathcal{L})$ takes values in continuous functions. Indeed, $\mathfrak{C}_{c}(r(U)\mathcal{G},\mathcal{L})$ is spanned by elements $f\in C_{c}(V,\mathcal{L})$ for some $V\in\operatorname{Bis}(\mathcal{G})$ with $r(V)\subseteq r(U)$, and if $\mathcal{G}$ is Hausdorff then $V\cap U$ is closed (and open) in $V$, and hence $f|_{U}$ is continuous for every $f\in C_{c}(V,\mathcal{L})$.
@@
Convolution defines a $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-bimodule structure on $\mathcal{B}(\mathcal{G},\mathcal{L})$ which is continuous in the sense that if $\{f_{n}\}_{n=1}^{\infty}\subseteq\mathcal{B}(\mathcal{G},\mathcal{L})$ converges to $f\in\mathcal{B}(\mathcal{G},\mathcal{L})$ in $\|\cdot\|_{\infty}$, then for any $g\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ we have $g*f_{n}\to g*f$ and $f_{n}*g\to f*g$ in $\|\cdot\|_{\infty}$. It restricts to a continuous $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-bimodule structure on $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$.
@@
Fix $g\in C_{c}(U^{*},\mathcal{L})$, where $U\in\operatorname{Bis}(\mathcal{G})$. For any $f\in\mathcal{B}(\mathcal{G},\mathcal{L})$ we have

### v3

@@ Acknowledgements
The research of KB and BK was supported by the National Science Center (NCN) grant no. 2019/35/B/ST1/02684. BK was partially supported by the National Science Centre, Poland, through the WEAVE-UNISONO grant no. 2023/05/Y/ST1/00046. BK thanks Agnieszka Stocka whose comments helped formulate the definition of reduced algebras in a correct way. We all thank the anonymous referees for pointing us to [Dix61] and [Elk25], and for helpful suggestions which improved the manuscript.
@@ 2. Preliminaries
We recall here some notation and facts from [BKM25]. Throughout this paper $\mathcal{G}$ denotes an étale groupoid with locally compact Hausdorff unit space $X$. Thus the range and domain maps $r,d:\mathcal{G}\to X$ are local homeomorphisms, $X$ is an open subset of $\mathcal{G}$, and $\mathcal{G}$ is Hausdorff if and only if $X$ is also closed in $\mathcal{G}$. We consider spaces, and algebras, over the field $\mathbb{F}=\mathbb{R},\mathbb{C}$ of real or complex numbers. By an ideal in a Banach algebra, unless stated otherwise, we mean a closed and two-sided ideal.
@@ 2.1. Twisted groupoids
As in [BKM25, Section 4], we define twists of $\mathcal{G}$ as Fell line bundles. A line bundle $\mathcal{L}$ over $\mathcal{G}$ is a topological space $\mathcal{L}=\bigsqcup_{\gamma\in\mathcal{G}}L_{\gamma}$ such that the canonical map $\mathcal{L}\to\mathcal{G}$ is continuous and open, and each fiber $L_{\gamma}\cong\mathbb{F}$ is a one dimensional Banach space with a structure consistent with the topology of $\mathcal{L}$ in the sense that the maps $\bigsqcup_{\gamma\in\mathcal{G}}L_{\gamma}\times L_{\gamma}\ni(z,w)\mapsto z+w\in\mathcal{L}$, $\mathbb{F}\times\mathcal{L}\ni(\lambda,z)\mapsto\lambda z\in\mathcal{L}$ and $\mathcal{L}\ni z\mapsto|z|\in\mathbb{R}$ are continuous. For any open $U\subseteq\mathcal{G}$, we write $C_{c}(U,\mathcal{L})$, $C_{0}(U,\mathcal{L})$, $C(U,\mathcal{L})$ and $C_{u}(U,\mathcal{L})$ for the spaces of continuous sections of $\mathcal{L}|_{U}:=\bigsqcup_{\gamma\in U}L_{\gamma}$ that are compactly supported, vanishing at infinity, bounded and unitary, respectively. We use similar notation $\mathcal{B}_{c}(U,\mathcal{L})$, $\mathcal{B}_{0}(U,\mathcal{L})$, $\mathcal{B}(U,\mathcal{L})$ when continuous is replaced by Borel. In particular, $\mathcal{B}_{0}(U,\mathcal{L})$ consists of Borel sections $f:U\to\mathcal{L}$ such that for every $\varepsilon>0$ the set $\{\gamma\in U:|f(\gamma)|\geq\varepsilon\}$ is precompact. We consider these spaces equipped with the supremum norm $\|f\|_{\infty}=\sup_{\gamma\in U}|f(\gamma)|$. Every line bundle is locally trivial, which means that every point has a neighborhood $U\subseteq\mathcal{G}$ such that $C_{u}(U,\mathcal{L})\neq\emptyset$. Note that, for any $c\in C_{u}(U,\mathcal{L})$, the map $f\mapsto f\cdot c$ establishes isometric isomorphisms $C_{*}(U,\mathcal{L})\cong C_{*}(U)$ and $\mathcal{B}_{*}(U,\mathcal{L})\cong\mathcal{B}_{*}(U)$ where * stands for either $c$, 0 or an empty space.
@@
A twist over the groupoid $\mathcal{G}$ is a line bundle $\mathcal{L}=\bigsqcup_{\gamma\in\mathcal{G}}L_{\gamma}$ over $\mathcal{G}$ equipped with two continuous maps $\cdot:\mathcal{L}\times\mathcal{L}\supseteq\bigsqcup_{d(\gamma)=r(\eta)}L_{\gamma}\times L_{\eta}\to\mathcal{L}$ and $^{*}:\mathcal{L}\to\mathcal{L}$ called multiplication and involution, such that:
@@
    $\cdot$ restricts to bilinear maps $L_{\gamma}\times L_{\eta}\to L_{\gamma\eta}$, $(\gamma,\eta)\in\mathcal{G}^{(2)}$, that make the norm $|\cdot|$ multiplicative, and $\cdot$ is associative, that is, $(\mathcal{L},\cdot)$ is a semigroupoid;
@@
The trivial bundle $\mathcal{L}=\mathcal{G}\times\mathbb{F}$ with pointwise operations is the trivial twist. More generally, if $\sigma:\mathcal{G}^{(2)}\to\{z\in\mathbb{F}:|z|=1\}$ is a continuous normalised $2$-cocycle on $\mathcal{G}$, that is, $\sigma\big(r(\gamma),\gamma\big)=1=\sigma\big(\gamma,d(\gamma)\big)$ and $\sigma(\alpha,\beta)\sigma(\alpha\beta,\gamma)=\sigma(\beta,\gamma)\sigma(\alpha,\beta\gamma)$ for every composable triple $\alpha,\beta,\gamma\in\mathcal{G}$, then we may consider a twist $\mathcal{L}_{\sigma}$ given by the trivial bundle $\mathcal{G}\times\mathbb{F}$ with operations given by $(\alpha,w)\cdot(\beta,z):=\big(\alpha\beta,\sigma(\alpha,\beta)wz\big)$, $(\alpha,w)^{*}:=\big(\alpha^{-1},\overline{\sigma(\alpha^{-1},\alpha)w}\big).$ Every twist based on a trivial line bundle is of this form. In general, twists defined above are equivalent to Kumjian’s twist [Kum86] defined using extensions of $\mathcal{G}$.
The set of open bisections $\operatorname{Bis}(\mathcal{G}):=\{U\subseteq\mathcal{G}\text{ open}:\text{$r|_{U}$ and $d|_{U}$ are injective}\}$ forms a unital inverse semigroup with operations defined by
@@
where $U,V\in\operatorname{Bis}(\mathcal{G})$. The unit space $X$ is the unit in $\operatorname{Bis}(\mathcal{G})$. For any twist $\mathcal{L}$ over $\mathcal{G}$, the set
[…]
forms a unital inverse subsemigroup $S(\mathcal{L})\subseteq\operatorname{Bis}(\mathcal{G})$ which is wide in the sense that it covers $\mathcal{G}$ and the intersection of any two elements in $S(\mathcal{L})$ is a union of elements from $S(\mathcal{L})$, see [BKM25, Lemma 4.2].
@@ 2.2. Banach algebras associated to twisted groupoids
where $\cdot$ and the final * indicate the product and involution from $\mathcal{L}$. When the groupoid $\mathcal{G}$ is Hausdorff this reduces to the usual *-algebra $C_{c}(\mathcal{G},\mathcal{L})$. We associate a norm on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ with any unital inverse subsemigroup $S\subseteq\operatorname{Bis}(\mathcal{G})$ which covers $\mathcal{G}$. By [BKM25, Lemma 4.5], the formula
@@
The class of algebras $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ defined above consists of Banach algebra Hausdorff completions of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ in seminorms $\|\cdot\|_{\mathcal{R}}$ for which there exists a unital inverse subsemigroup $S\subseteq\operatorname{Bis}(\mathcal{G})$ covering $\mathcal{G}$ such that $\|\cdot\|_{\mathcal{R}}\leq\|\cdot\|_{\infty}$ on every $C_{c}(U,\mathcal{L})$, $U\in S$ (this is equivalent to $\|\cdot\|_{\mathcal{R}}\leq\|\cdot\|_{\textrm{max}}^{S}$ on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$). Without loss of generality, we may always assume here that $S$ is closed under taking subsets, so that in particular it is wide, and that $S\subseteq S(\mathcal{L})$, see [BKM25, Remark 4.7].
@@
Let $p\in[1,\infty]$. We recall (see e.g. [Phi13b]) that a Banach algebra $A$ is an $L^{p}$-operator algebra if there is an isometric representation $\psi:A\to B(L^{p}(\mu))$, where $L^{p}(\mu)$ is the Banach $L^{p}$-space associated to some measure $\mu$. Following [BKM25, Definition 5.12], we define the full $L^{p}$-operator algebra of $(\mathcal{G},\mathcal{L})$ as $F^{p}(\mathcal{G},\mathcal{L}):=F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ where $\mathcal{R}$ consists of representations $\psi:F^{S}(\mathcal{G},\mathcal{L})\to B(E)$, for some unital inverse subsemigroup $S\subseteq\operatorname{Bis}(\mathcal{G})$ which covers $\mathcal{G}$, such that
@@
    operators in $\psi(C_{c}(X)^{+})$ preserve the cone of positive functions in $E$.
The following comments hold by [BKM25, Theorem 5.13]. The above definition does not depend on the choice of $S$, and so $F^{p}(\mathcal{G},\mathcal{L})$ is $\operatorname{Bis}(\mathcal{G})$-graded. In fact, denoting by $\|\cdot\|_{L^{p}}$ the norm in $F^{p}(\mathcal{G},\mathcal{L})$, for $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ and $q\in[1,\infty]$ with $1/p+1/q=1$ (we put $q=1$ if $p=\infty$ and vice versa) we get
@@
in the complex case. In the real case trouble may come from representations as in [BKM25, Example 2.14]. For $\mathbb{F}=\mathbb{R},\mathbb{C}$ and $p=2$, $\|\cdot\|_{L^{2}}$ is the maximal submultiplicative norm on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ satisfying the $C^{*}$-equality. Thus $C^{*}(\mathcal{G},\mathcal{L}):=F^{2}(\mathcal{G},\mathcal{L})$ is a (real or complex) $C^{*}$-algebra, which is universal in the sense that *-homomorphisms from $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ into a $C^{*}$-algebra $B$ extend to representations of $C^{*}(\mathcal{G},\mathcal{L})$ in $B$, and every representation of $C^{*}(\mathcal{G},\mathcal{L})$ in $B$ is automatically *-preserving. The norm $\|\cdot\|_{L^{\infty}}$ does not change, if we replace the $L^{\infty}$-space in (R1) by $E=C_{0}(\Omega)$ for some locally compact Hausdorff space. Thus $F^{\infty}(\mathcal{G},\mathcal{L})$ could also be viewed as the full $C_{0}$-operator algebra of $(\mathcal{G},\mathcal{L})$. Moreover, if $q\in[1,\infty]$ satisfies $1/p+1/q=1$, then the involution in $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ extends to an isometric anti-isomorphism $F^{p}(\mathcal{G},\mathcal{L})\stackrel{{\scriptstyle\mathrm{anti}}}{{\cong}}F^{q}(\mathcal{G},\mathcal{L})$.
@@
defines a representation $\Lambda_{p}:F^{p}(\mathcal{G},\mathcal{L})\to B(\ell^{p}(\mathcal{G},\mathcal{L}))$ which is injective on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, see [BKM25, Proposition 5.1]. We call $F_{\operatorname{r}}^{p}(\mathcal{G},\mathcal{L}):=F_{\operatorname{r}}^{\{\Lambda_{p}\}}(\mathcal{G},\mathcal{L})$ the reduced $L^{p}$-operator algebra of $(\mathcal{G},\mathcal{L})$ and we denote the norm by $\|\cdot\|_{L^{p},\operatorname{r}}$. By [BKM25, Theorem 5.13] we have
@@ 3. Reduced groupoid Banach algebras
We fix a twisted groupoid $(\mathcal{G},\mathcal{L})$. In this section, we will consider completions of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ that are equipped with ‘generalized conditional expectations’ that extend the map $E_{X}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to\mathcal{B}_{c}(X)$, given by restriction $f\mapsto f|_{X}$. When $\mathcal{G}$ is Hausdorff, then $E_{X}:C_{c}(\mathcal{G},\mathcal{L})\to C_{c}(X)\subseteq C_{c}(\mathcal{G},\mathcal{L})$ is a projection.
@@
For any unital inverse subsemigroup $S\subseteq\operatorname{Bis}(\mathcal{G})$ covering $\mathcal{G}$, the algebra $F^{S}(\mathcal{G},\mathcal{L})$ is a groupoid Banach algebra. For any $p\in[1,\infty]$, the $L^{p}$-operator algebras $F^{p}(\mathcal{G},\mathcal{L})$ and $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ are groupoid Banach algebras and $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ is reduced, cf. Proposition 3.15 below. There is a unique $C^{*}$-norm $\|\cdot\|_{C^{*}_{\operatorname{r}}}$ on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ yielding a reduced Banach algebra, and then this algebra coincides with the standard reduced $C^{*}$-algebra $C^{*}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ of $(\mathcal{G},\mathcal{L})$ (or its real counterpart). The full $C^{*}$-algebra $C^{*}(\mathcal{G},\mathcal{L})$, which is a completion of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ in the maximal $C^{*}$-norm $\|\cdot\|_{C^{*}\max}$, is a groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$. Thus a $C^{*}$-norm $\|\cdot\|_{\mathcal{R}}$ on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ yields a groupoid Banach algebra if and only if it satisfies $\|\cdot\|_{C^{*}_{\operatorname{r}}}\leq\|\cdot\|_{\mathcal{R}}\leq\|\cdot\|_{C^{*}\max}$.
We make some preparations to reveal more structure in groupoid Banach algebras. For any $U\in\operatorname{Bis}(\mathcal{G})$, we view the space of bounded Borel sections $\mathcal{B}(U,\mathcal{L})$ as a Banach $C(X)$-bimodule:
@@
gives a $C_{0}(X)$-bilinear map $E_{U}:\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to\mathcal{B}(U,\mathcal{L})$. The map $E_{U}$ on $\mathfrak{C}_{c}(r(U)\mathcal{G},\mathcal{L})=C_{c}(r(U))*\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, takes values in compactly supported Borel sections $\mathcal{B}_{c}(U,\mathcal{L})=C_{c}(r(U))*\mathcal{B}(U,\mathcal{L})$. When $\mathcal{G}$ is Hausdorff $E_{U}:\mathfrak{C}_{c}(r(U)\mathcal{G},\mathcal{L})\to C_{c}(U,\mathcal{L})$ takes values in continuous functions. Indeed, $\mathfrak{C}_{c}(r(U)\mathcal{G},\mathcal{L})$ is spanned by elements $f\in C_{c}(V,\mathcal{L})$ for some $V\in\operatorname{Bis}(\mathcal{G})$ with $r(V)\subseteq r(U)$, and if $\mathcal{G}$ is Hausdorff then $V\cap U$ is closed (and open) in $V$, and hence $f|_{U}$ is continuous for every $f\in C_{c}(V,\mathcal{L})$.
@@
Convolution defines a $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-bimodule structure on $\mathcal{B}(\mathcal{G},\mathcal{L})$ which is continuous in the sense that if a sequence $(f_{n})_{n=1}^{\infty}$ in $\mathcal{B}(\mathcal{G},\mathcal{L})$ converges to $f\in\mathcal{B}(\mathcal{G},\mathcal{L})$ in $\|\cdot\|_{\infty}$, then, for any $g\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, we have $g*f_{n}\to g*f$ and $f_{n}*g\to f*g$ in $\|\cdot\|_{\infty}$. It restricts to a continuous $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-bimodule structure on $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$.
@@
Fix $g\in C_{c}(U^{*},\mathcal{L})$, where $U\in\operatorname{Bis}(\mathcal{G})$. For any $f\in\mathcal{B}(\mathcal{G},\mathcal{L})$, we have